l. Air flows steadily in a pipe with a velocity of 60 ft/s. Surrounding the pipe in an annulus is a second flow of air with a velocity of 40 ft/s. Both flows are exhausted through a 1.5 ft dia pipe. If the velocity is uniform at the exit, determine the velocity at the exit. Assume constant density. 2 ft dia exit 1 ft dia
Answer #1

According to equation of continuity, (mass flown per second) \SigmaQin = \SigmaQout

This can be considered as a flow system with two inlets and one exit.

Then, \SigmaQin   = A1v1\rho + A2 v2\rho      and \SigmaQout = A3v3\rho

As \rho is the same, we have: A1v1 + A2 v2 = A3v3 ..................(i)

Here, take d1 = 1 ft, d2 = 2 ft and d3 = 1.5 ft. Also, v1 = 60 ft/s, v2 = 40 ft/s, v3 = ?

A1 = \pi(d12 )/4, A2 = \pi[(d22 – d12)/4] and A3 = \pi(d32)/4

[ Note: d stands for diameter.. so area = \pid2/4. A2 is the annular area. hence the formula].

Putting A1, A2 and A3 in equation (i),

\pi(d12 /4) × v1 + \pi[(d22 – d12)/4] × v2 = \pi(d32)/4 × v3

Equation reduces to

d12 × v1 + (d22 – d12) × v2 = d32 × v3  

    ==> 1 × 60 + (4 – 1) × 40 = 1.52 × v3

    ==> 2.25 v3 = 180 ==> v3 = 80 ft/s.....(Required Ans).

Hope it is clear..